# Exploring the operational limits of Mirnov coil arrays in stellarators by means of a synthetic diagnostic

### v1

@@ Abstract
A synthetic Mirnov coils diagnostic is presented and used to study the capabilities of the poloidal array of single-axis coils and the two helical arrays of tri-axial coils installed in the TJ-II stellarator. This tool integrates the plasma currents induced by Alfvén-like perturbations of the electric potential inside the plasma and provides the induced magnetic field oscillations anywhere outside of it. The simulated signals can then be analyzed in the same manner as the experimental ones, and a scan on the radial position and width of the potential perturbation is conducted to find the limiting values that produce identifiable signals. We find, not surprisingly, that core-localized modes are indistinguishable from one another; and that the identification of low-n, low-m modes is often subject to off-by-one errors. We also determine the optimal polarization basis in which to analyze the tri-axial coils signals and address the diagnostic performance when resolving components of gap modes such as HAEs. Additionally, selected cases have been analyzed with a simplified plasma response model, showing that plasma shielding of the mode currents may further deteriorate the accuracy of the mode identification method. The code for this synthetic diagnostic, that can be easily used to conduct similar analysis for other devices, is publicly available in the following GitHub repository: https://github.com/pponsv/synth_mirnov.
@@ I Introduction
The experimental determination of the spatial structure of AEs through the identification of its poloidal (m) and toroidal (n) mode numbers is an integral part of this task, as it allows for theory-experiment comparisons. Of the set of diagnostics that provide information on mode numbers, Mirnov coil arrays distributed around the plasma columnMirnov (1965) are one of the more widely used. Typically, tokamaks currently in operation like DIII-D Strait (2006), ASDEX Upgrade Horváth et al. (2016); Schittenhelm and Zohm (1997) or TCV Moret et al. (1998) have a large number of poloidal Mirnov coils and a much smaller number of toroidal ones, that are usually enough for low-$n$ modesMink et al. (2016). An exception to this trend is MAST-U and its OMAHA poloidal array Hole et al. (2009) with only 10 coils that are optimized to minimize the spatial aliasing of the measured signal and thus are able to resolve high mode numbers with a comparatively sparse arrangement of sensors. Thus, poloidal mode numbers of most common AEs are usually well identified in tokamaks.
@@
The physics underlying the AEs model used in the synthetic diagnostic and the plasma response calculation is described in appendix A. Section II describes the spatial distribution of the different sets of Mirnov coils installed inside the device. One of them, which measures only one component of the field at each position, is indicated for the measurement of the poloidal mode number while the other two are helical sets of triaxial coils measuring the three components of the magnetic field at each position. Section III describes the mode number analysis technique applied to both synthetic and real signals, and section IV describes the results of the simulations and addresses the performance of the arrays taking into account aspects such as mode position in the plasma or magnetic field polarization at the measurement locations, which is related to mode polarization in the plasma in a non-trivial way due to the highly three-dimensional structure of the equilibrium field. In section V, the results of the simulations including the plasma response model are presented.
@@ II Experimental set-up
Figure 2 depicts the design orientations of the coils. By construction, the uncertainty in the orientation of the poloidal array coils is low, as they are perfectly aligned inside a rigid metallic tube. The tri-axial sensors, on the other hand, are located inside a semi-rigid corrugated tube that had to be twisted for its installation inside the vacuum vessel. The alignment of each tri-axial set of coils inside the tube is more sensitive to positioning errors and a calibration was necessary to determine their true orientations. Once these orientations are known, the signal can be projected over arbitrary directions in software, as the coils in each sensor are orthogonal. A comprehensive description of the helical array, its characteristics and the related calibration experiments can be found in ref. 21. Since this is a numerical study of the performance of the diagnostic, we will assume no coil orientation errors.
@@ III Mode number analysis
The Lomb-Scargle periodogram Lomb (1976); Scargle (1982) is widely used in astrophysics to identify periodicities in non-evenly-spaced observation data. In plasma physics, a 3D generalization was introduced by ZegenhagenZegenhagen et al. (2006) to conduct mode analysis with non-equispaced coils in W7-AS. The Lomb-Scargle periodogram fits the data $y_{ij}$ to a sinusoidal model:
$y_{ij}+\epsilon_{ij}=a\cos(\alpha_{ij})+b\sin(\alpha_{ij})$ (1)
@@
The periodogram is given by:
$P(\omega,n,m)=\dfrac{1}{YY}\left(\dfrac{\left[\sum_{ij}y_{ij}\cos(\alpha_{ij}-\tau)\right]^{2}}{\sum_{ij}\cos^{2}(\alpha_{ij}-\tau)}+\dfrac{\left[\sum_{ij}y_{ij}\sin(\alpha_{ij}-\tau)\right]^{2}}{\sum_{ij}\sin^{2}(\alpha_{ij}-\tau)}\right)$ (2)
@@
For a more in-depth discussion on this phase shift term, that differs from the one given in Zegenhagen et al. (2006), see appendix B. With this definition, the periodogram is normalized, giving results in the range $[0,1]$, with $P=0$ corresponding to zero agreement with the measured signals (simulations in this case) and $P=1$ corresponding to perfect agreement. When applied to experimental data, the analysis starts by finding the frequency $f_{0}$ of the detected eigenmode from an spectrogram obtained either via the Short-Time Fast Fourier Transform (STFFT) or the DMUSIC (Damped Multiple Signal Classification) method Kleiber et al. (2021). At this time, the frequency is selected manually, but mode following algorithms Vaz Mendes et al. (2023); Heinrich et al. (2024) can be readily applied for an automated analysis once the method has been shown to be sound. Then, a scan is performed over the relevant $n$ and $m$ mode numbers, that is, the ones that can be resolved given the distance between coils. The results are plotted as a 2D colormap of $P(n,m)$. For our present purposes, mode frequency and mode numbers are inputs for the synthetic modelling and the potential perturbation associated with the mode is given by equations 6 and 7.
@@
the measured signals therefore depend both on the toroidal and the poloidal mode numbers. This link between poloidal and toroidal angles limits the identifiable mode numbers, introducing false positives in a phenomenon akin to aliasing. The phase difference between signals measured by two coils separated $\Delta\varphi$ in the toroidal direction will be given by:
$\delta=(n-N_{\text{fp}}m)\Delta\varphi=l\Delta\varphi$ (5)
@@
The periodograms of each sub-array individually are virtually identical, a stripe with slope $1/N_{\text{fp}}$, but the extra poloidal information given by the sub-array separation in the poloidal direction is enough to partially remove the uncertainty. Only after adding the measurements of the poloidal array can we arrive at an unambiguous determination of both toroidal and poloidal mode numbers. This can be seen best in the bottom row of figure 4, that shows Lomb periodograms for the poloidal array (PA panel), the two helical arrays, combined (HA panel), and the product of the previous two. There, the dependency on the toroidal angle $\varphi$ of the poloidal array shows up as a slight slope in the mode identification of the PA panel, that otherwise (if the array had $\varphi=\text{constant}$) would appear as a horizontal stripe. Similarly, in the HA panel, a diagonal stripe appears, with slope $1/N_{\text{fp}}$ as discussed before. Having both sub-arrays provides enough information to discard some of the modes, and that is the reason why the stripe is not continuous. The most robust way to perform the mode identification given these limitations in the coil arrangement has proved to be taking the product of the PA and HA periodograms and finding its maximum.
@@ IV Simulations
Similarly to what happens in tokamaks, where modes of the same $n$ and different $m$ can evolve jointly at the same frequency within the TAE gap, the existence of helicity gaps (HAE) in stellarators allows for complex radially extended and weakly damped structures characterized by modes with different values of $n$ and $m$. For instance, in the case of an HAEμν gap created by the appropriate components in the Fourier expansions of the magnetic field and the metric coefficients, modes with $n_{1}$ and $m_{1}$ can couple with modes with $n_{2}=n_{1}+\nu N$ and $m_{2}=m_{1}+\mu$. Therefore, to account for arbitrary mode coupling, the synthetic diagnostic considers potential perturbations (see appendix A) of the form:
$\delta\phi(s,\theta,\varphi,t)=\sum_{mn}\delta\phi_{mn}^{\omega}(s)e^{i(m\theta+n\varphi-\omega t)}$ (6)
@@ IV.1 Single mode simulations
We start by running the synthetic diagnostic code for single mode perturbations defined by their mode numbers $n$, $m$ and their frequency $\omega$. The spatial structure of the electric potential perturbation is given by equation 6, and the radial profile $\delta\phi_{mn}^{\omega}$ is taken to be Gaussian:
$\delta\phi_{mn}^{\omega}(s)=A_{mn}^{\omega}\exp\left[-\dfrac{(s-s_{0})^{2}}{\sigma}\right]$ (7)
with $A_{mn}^{\omega}\in\mathbb{C}$. In figure 5 the potential perturbation for a single $m=5$, $n=7$, $s_{0}=0.3$, $\sigma=0.015$ mode is shown for several toroidal cuts at constant $\varphi$ in a half period of the device. The current is calculated from the vector potential perturbation using equation 22, which we rewrite here for convenience:
@@
Figure 6 shows the parallel component of the vector potential and the component along $\mathbf{e}_{\varphi}$ of the induced current associated to the potential pertubation shown in figure 5. The structure of the current perturbation is notably more complex than that of its potential, which is to be expected since it arises from the double curl of the vector potential. Also, from this expression we see that narrower modes will have higher current densities, as the slopes will be more steep and thus the derivatives will be bigger.
Figure 6: 2D projection of the plasma cross-section at constant toroidal boozer angle $\varphi=\pi/4$, showing the parallel component of the perturbed vector potential ($\delta A_{\parallel}$) in (a) and the contravariant $\varphi$ component of the current perturbation ($\delta J_{\text{ext}}^{\varphi}$) in (b), for the potential perturbation shown in figure 5.
@@ IV.1.1 Poloidal array
We begin by evaluating the ability of the poloidal array to identify poloidal mode numbers. For that, we must take into account that the signal measured by the probes is $\delta\dot{B}_{\text{p}}=\delta\dot{\mathbf{B}}\cdot\mathbf{P}$, where $\mathbf{P}$ is the vector perpendicular to the coil winding plane (see figure 2), and lies in the RZ plane. Even though the physical diagnostic only measures $\delta\dot{\mathbf{B}}$ along a predetermined direction, the synthetic diagnostic provides the full 3D signal, opening up the possibility of an optimization study on the best possible orientation of a set of single axis coils for mode identification. This is treated at the end of this section.
@@
This outcome must be taken into account when analising experimental data since the identification of the core-localized mode number invariably fails. This is not a limitation of the analysis technique, but rather a well-known limitation of this type of diagnostic that cannot be avoided. The structure of core-localized modes cannot be experimentally characterized by measurements in the periphery, as is the case for magnetic diagnostics. Care must be taken into not trying to analise a mode outside of the diagnostic’s operational limits. For that, diagnostics that can provide potential perturbation profiles, such as HIBP Melnikov et al. (2022) or SXR tomographyDreval et al. (2021), are the ideal candidate to complement the analysis.
To map these operational limits, a series of scans in both poloidal and toroidal mode number, perturbation width, and radial location have been carried out. The pattern of core modes losing phase information is common, as we can see in figure 9, where the difference between the simulated mode number ($m_{\text{in}}$) and the measured mode number ($m_{\text{out}}$) is represented. In this figure, we can also see that the performance for low-m modes is slightly worse than for higher-m modes, presenting phase information loss for modes further out from the core and more off-by-one errors. The poloidal array does not cover the full 360 degrees because the plasma column is very close to the vacuum chamber in the zone of highest indentation of the magnetic surfaces, and it is this lack of coverage that reduces its accuracy for low-m modes. When studying higher-m modes, as in figure 10, the performance remains reasonably good for very high mode numbers, even above the nyquist mode number
$m_{\text{nyq}}\simeq\dfrac{1}{2}\dfrac{360\cdot 25}{270}\simeq 16$ (9)
@@
This is due to the ability of the Lomb periodogram to take advantage of the slightly uneven spacing of the coils in boozer coordinates VanderPlas (2018). For narrower modes, however, the phase information loss appears for more outer modes beginning at $m\sim 14$.
Now we turn to study the polarization of the detected perturbation. Inside the plasma, the transverse polarization of shear Alfvén waves makes the perturbed magnetic field perpendicular to the equilibrium field; but this does not hold outside, where the field perturbation detected at a given spatial position is given by an integral over the whole plasma volume. Finding the right polarization base to guarantee an optimal result is an essential part of the analysis.

### v3

# thanks: J.A. Alonso et al 2024 Nucl. Fusion 64 112018

@@ Abstract
A synthetic Mirnov coils diagnostic for non-axisymmetric magnetic configurations is presented and used to study the capabilities of the poloidal array of single-axis coils and the two helical arrays of tri-axial coils installed in the TJ-II stellarator. This tool integrates the plasma currents induced by Alfvén-like perturbations of the electric potential inside the plasma and provides the induced magnetic field oscillations anywhere outside of it. The simulated signals can then be analyzed in the same manner as the experimental ones, and a scan on the radial position and width of the potential perturbation is conducted to find the limiting values that produce identifiable signals. We find, not surprisingly, that core-localized modes are indistinguishable from one another; and that the identification of low-n, low-m modes is often subject to off-by-one errors. We also determine the optimal polarization basis in which to analyze the tri-axial coils signals and address the diagnostic performance when resolving components of gap modes. Additionally, selected cases have been analyzed with a simplified plasma response model, showing that plasma shielding of the mode currents may further deteriorate the accuracy of the mode identification method. We conclude with the analysis of an experimental case taken from the TJ-II database to illustrate the usefulness of the diagnostic.
@@ I Introduction
The experimental determination of the spatial structure of AEs through the identification of its poloidal (m) and toroidal (n) mode numbers is an integral part of this task, as it allows for theory-experiment comparisons. Of the set of diagnostics that provide information on mode numbers, Mirnov coil arrays distributed around the plasma columnMirnov (1965) are one of the more widely used. Typically, tokamaks currently in operation like DIII-D Strait (2006), ASDEX Upgrade Horváth et al. (2016); Schittenhelm and Zohm (1997) or TCV Moret et al. (1998) have a large number of poloidal Mirnov coils and a much smaller number of toroidal ones, that are usually enough for low-$n$ modesMink et al. (2016). An exception to this trend is MAST-U and its OMAHA poloidal array Hole et al. (2009) with only 10 coils that are optimized to minimize the spatial aliasing of the measured signal and thus are able to resolve high mode numbers with a comparatively sparse arrangement of sensors. Compared to stellarators, the toroidal and poloidal mode numbers of the most common modes are more easily identified in tokamaks, with the poloidal number being the most problematic.
@@
The physics underlying the AEs model used in the synthetic diagnostic and the plasma response calculation is described in appendix A. Section II describes the spatial distribution of the different sets of Mirnov coils installed inside the device. One of them, which measures only one component of the field at each position, is indicated for the measurement of the poloidal mode number while the other two are helical sets of triaxial coils measuring the three components of the magnetic field at each position. Section III describes the mode number analysis technique applied to both synthetic and real signals, and section IV describes the results of the simulations and addresses the performance of the arrays taking into account aspects such as mode position in the plasma or magnetic field polarization at the measurement locations, which is related to mode polarization in the plasma in a non-trivial way due to the highly three-dimensional structure of the equilibrium field. Section V showcases the analysis of a discharge of TJ-II, where the lessons learned are applied to experimental data. Our conclusions and further discussion are held in section VI.
@@ II Experimental set-up
Figure 2 depicts the design orientations of the coils. By construction, the uncertainty in the orientation of the poloidal array coils is low, as they are perfectly aligned inside a rigid metallic tube. The tri-axial sensors, on the other hand, are located inside a semi-rigid corrugated tube that had to be twisted for its installation inside the vacuum vessel. The alignment of each tri-axial set of coils inside the tube is more sensitive to positioning errors and a calibration was necessary to determine their true orientations. Once these orientations are known, the signal can be projected over arbitrary directions in software, as the coils in each sensor are orthogonal. A comprehensive description of the helical array, its characteristics and the related calibration experiments can be found in ref. 21. For the numerical study of the diagnostic performance we have assumed ideal coils distributed according to the position and orientation of the original design, thus avoiding adding an extra layer of complexity. However, the calibration is essential for the analysis of the experimental data and has been taken into account in the analysis presented in section V.
@@ III Mode number analysis
The Lomb-Scargle periodogram Lomb (1976); Scargle (1982) is widely used in astrophysics to identify periodicities in non-evenly-spaced observation data. In plasma physics, a 3D generalization was introduced by ZegenhagenZegenhagen et al. (2006) to conduct mode analysis with non-equispaced coils in W7-AS. The Lomb-Scargle periodogram fits the data $y_{ij}$ to the sinusoidal model
$y_{ij}+\epsilon_{ij}=a\cos(\alpha_{ij})+b\sin(\alpha_{ij})\,,$ (1)
@@
The periodogram is given by
$P_{\text{LS}}(\omega,n,m)=\dfrac{1}{YY}\left(\dfrac{\left[\sum_{ij}y_{ij}\cos(\alpha_{ij}-\tau)\right]^{2}}{\sum_{ij}\cos^{2}(\alpha_{ij}-\tau)}+\dfrac{\left[\sum_{ij}y_{ij}\sin(\alpha_{ij}-\tau)\right]^{2}}{\sum_{ij}\sin^{2}(\alpha_{ij}-\tau)}\right)\,,$ (2)
@@
$\tan 2\tau=\dfrac{\sum_{ij}\sin 2\alpha_{ij}}{\sum_{ij}\cos 2\alpha_{ij}}\,.$ (3)
For a more in-depth discussion on this phase shift term, that differs from the one given in Zegenhagen et al. (2006), see appendix B. With this definition, the periodogram is normalized, giving results in the range $[0,1]$, with $P_{\text{LS}}=0$ corresponding to zero agreement with the measured signals (simulations in this case) and $P_{\text{LS}}=1$ corresponding to perfect agreement. When applied to experimental data, the analysis starts by finding the frequency $f_{0}$ of the detected eigenmode from an spectrogram obtained either via the Short-Time Fast Fourier Transform (STFFT) or the DMUSIC (Damped Multiple Signal Classification) method Kleiber et al. (2021). At this time, the frequency is selected manually, but mode following algorithms Vaz Mendes et al. (2023); Heinrich et al. (2024) can be readily applied for an automated analysis once the method has been shown to be sound. Then, a scan is performed over the relevant $n$ and $m$ mode numbers, that is, the ones that can be resolved given the distance between coils. The results are plotted as a 2D colormap of $P_{\text{LS}}(n,m)$. For our present purposes, mode frequency and mode numbers are inputs for the synthetic modelling and the potential perturbation associated with the mode is given by equations 6 and 7.
@@
the measured signals therefore depend both on the toroidal and the poloidal mode numbers. This link between poloidal and toroidal angles limits the identifiable mode numbers, introducing false positives in a phenomenon akin to aliasing. The phase difference between signals measured by two coils separated $\Delta\varphi$ in the toroidal direction will be:
$\delta=(n-N_{\text{fp}}m)\Delta\varphi=l\Delta\varphi\,,$ (5)
@@
The periodograms of each sub-array individually are virtually identical, a stripe with slope $1/N_{\text{fp}}$, but the extra poloidal information given by the sub-array separation in the poloidal direction is enough to partially remove the uncertainty. Only after adding the measurements of the poloidal array can we arrive at an unambiguous determination of both toroidal and poloidal mode numbers. This can be seen best in the bottom row of figure 4, that shows Lomb periodograms for the poloidal array (PA panel), the two helical arrays, combined (HA panel), and the product of the previous two. There, the dependency on the toroidal angle $\varphi$ of the poloidal array shows up as a slight slope in the mode identification of the PA panel, that otherwise (if the array had $\varphi=\text{constant}$) would appear as a horizontal stripe. Similarly, in the HA panel, a diagonal stripe appears, with slope $1/N_{\text{fp}}$ as discussed before. Having both sub-arrays provides enough information to discard some of the modes, and that is the reason why the stripe is not continuous. The most robust way to perform the mode identification given these limitations in the coil arrangement has proved to be taking the product of the PA and HA periodograms and finding its maximum. To our knowledge, this is the first time that this method has been used.
@@ IV Simulations
We divide this section into two main subsections, one of them dedicated to the simulations with a single synthetic mode and the other to the simulations dealing with coupled modes. In the first subsection the essential aspects to be taken into account are first illustrated using the poloidal array. We discuss the treatment of the mode polarization, that is necessary for the mode analysis using the helical arrays. In the second subsection, where we study coupled mode simulations with both arrays, our goal is to investigate the ability of the diagnostic to resolve different pairs of physically relevant coupled modes. To account for arbitrary mode coupling, the synthetic diagnostic considers potential perturbations (see appendix A) of the form
$\delta\phi(s,\theta,\varphi,t)=\sum_{mn}\delta\phi_{mn}^{\omega}(s)e^{i(m\theta+n\varphi-\omega t)}\,,$ (6)
@@
The spatial structure of the electrostatic potential perturbation is given by equation 6, and the radial profile $\delta\phi_{mn}^{\omega}$ is taken to be Gaussian:
$\delta\phi_{mn}^{\omega}(s)=A_{mn}^{\omega}\exp\left[-\dfrac{(s-s_{0})^{2}}{\sigma}\right]\,,$ (7)
with $A_{mn}^{\omega}\in\mathbb{C}$. The current is calculated from the vector potential perturbation using equation 22, which we rewrite here for convenience:
@@ IV.1 Single mode simulations
We start by running the synthetic diagnostic code for single mode perturbations defined by their mode numbers $n$, $m$ and their frequency $\omega$. In figure 5 the potential perturbation for a single $m=5$, $n=7$, $s_{0}=0.3$, $\sigma=0.015$ mode is shown for several toroidal cuts at constant $\varphi$ in a half period of the device.
@@
Figure 6: 2D projection of the plasma cross-section at constant toroidal boozer angle $\varphi=\pi/4$, showing the parallel component of the perturbed magnetic vector potential ($\delta A_{\parallel}$) in (a) and the contravariant $\varphi$ component of the current perturbation ($\delta J_{\text{ext}}^{\varphi}$) in (b), for the potential perturbation shown in figure 5.
Figure 6 shows the parallel component of the vector potential and the component along $\mathbf{e}_{\varphi}$ of the induced current associated to the potential pertubation shown in figure 5. The structure of the current perturbation is notably more complex than that of its potential, which is to be expected since it arises from the double curl of the vector potential. Also, from this expression we see that narrower modes will have higher current densities, as the slopes will be more steep and thus the derivatives will be larger.
@@ IV.1.1 Poloidal array
We begin by evaluating the ability of the poloidal array to identify poloidal mode numbers. For that, we must take into account that the signal measured by the probes is $\delta\dot{B}_{\text{p}}=\delta\dot{\mathbf{B}}\cdot\mathbf{P}$, where $\mathbf{P}$ is the vector perpendicular to the coil winding plane (see figure 2), and lies in the RZ plane.
@@
This outcome must be taken into account when analising experimental data since the identification of the core-localized mode number invariably fails. This is not a limitation of the analysis technique, but rather a well-known limitation of this type of diagnostic that cannot be avoided. The structure of core-localized modes cannot be experimentally characterized by measurements in the periphery, as is the case for magnetic diagnostics. For instance, observations in tokamaks Sharapov et al. (2004) have shown that magnetic fluctuation diagnostics are indeed less sensitive to core-localized modes and therefore, care must be taken into not trying to analise a mode outside of the diagnostic’s operational limits. For that, diagnostics such as HIBP Melnikov et al. (2022), that can provide potential perturbation profiles, or tomography of soft X-ray (SXR) measurements, proportional to the plasma fluctuations structure Dreval et al. (2021), are the ideal candidate to complement the analysis.
@@
To map these operational limits, a series of scans in both poloidal and toroidal mode number, perturbation width, and radial location have been carried out. The pattern of core modes losing phase information is common, as we can see in figure 9, where the difference between the simulated mode number ($m_{\text{in}}$) and the measured mode number ($m_{\text{out}}$) is represented. In this figure, we can also see that the performance for low-m modes is slightly worse than for higher-m modes, presenting phase information loss for modes further out from the core and more off-by-one errors. The poloidal array (figure 1, green) does not cover the full 360 degrees because the plasma column is very close to the vacuum chamber in the zone of highest indentation of the magnetic surfaces, and it is this lack of coverage that reduces its accuracy for low-m modes. When studying higher-m modes, as in figure 10, the performance remains reasonably good for very high mode numbers, even above the nyquist mode number
$m_{\text{nyq}}\simeq\dfrac{1}{2}\dfrac{360\cdot 25}{270}\simeq 16\,.$ (9)
This is due to the ability of the Lomb periodogram to take advantage of the slightly uneven spacing of the coils in boozer coordinates VanderPlas (2018). For narrower modes, however, the phase information loss appears for more outer modes beginning at $m\sim 14$. High $n$ mode numbers have been used in figure 10 because the simultaneous detection of high mode numbers in both the poloidal and toroidal direction should be the most challenging situation for the diagnostic, and so that a meaningful comparison of the performance after the addition of the helical array in section IV.1.3 can be carried out. We will see later that this is not the case, and high mode numbers can be resolved remarkably well.
IV.1.2 Wave polarization
Now we turn to study the polarization of the detected perturbation. Wave polarization is a property of the wave that specifies the spatial orientation of the electric and magnetic field as the wave evolves in time. It carries information on the amplitude of the different components of the wave and the phase shifts between them in a given basis of polarization vectors Jackson (1999). The locus of the $\delta\mathbf{B}$ vector during one wave period is known as the polarization ellipse and lies in a plane whose orientation in space depends precisely on the polarization properties of the wave. Inside the plasma, the transverse polarization of shear Alfvén waves makes the perturbed magnetic field perpendicular to the equilibrium field; but this does not hold outside, where the field perturbation detected at a given spatial position is given by an integral over the whole plasma volume. Even though the poloidal array only measures $\delta\dot{\mathbf{B}}$ along a predetermined direction, the synthetic diagnostic provides the full 3D signal, opening up the possibility of an optimization study on the best possible orientation of a set of single axis coils for mode identification. Finding the right polarization base to guarantee an optimal result is an essential part of the analysis.